Operation Summary for December 13, 2010 - December 15, 2010
Monday Dec. 13
The day shift began with Operations monitoring the Tevatron (TeV) store 8361 (180.15E30) and the Recycler stash (123.2E10), with the Antiproton Source (Pbar) stacking (19.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-994 (JASMIN) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the evening shift, at 4:35 PM, store 8361 aborted due to a CIA (vacuum) crate failure at TeV sector F1. Operations called in Controls personnel to make repairs. Operators had the TeV ready for HEP shot setup by the end of the shift. Operations reported that the MI Power Supply Interlock Loop dropped out multiple times at MI-20 due to a flaky power supply door interlock switch; Operations called in EE Support personnel to repair the switch.
Tuesday Dec. 14
The midnight shift began with Operations monitoring the stash (428.2E10), with Pbar stacking (15.1E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations put the TeV into shot setup and established store 8363 at 1:30 AM with an average initial luminosity (AIL) of 312E30. Operations reported that the F1 back-up air compressor tripped off; operators managed to recover the main house compressor before the vacuum valves closed.
During the day shift, Operations terminated store 8363 at 1:20 PM and put the TeV into shot setup; Operations reported a short delay during the shot setup for wet engine maintenance at A3 and C1. Operations established store 8364 at 3:15 PM with an AIL of 339E30. Efforts to improve the MI tune continued throughout the day shift.
During the evening shift, Operations called in Booster personnel to replace a power supply regulator (B:VL4). Operators continued tuning MI to improve efficiency.
Wednesday Dec. 15
The midnight shift began with Operations monitoring store 8364 (114.15E30) and the stash (239.7E10), with Pbar stacking (14.8E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8364 at 6:50 AM and put the TeV into shot setup. Operations attempted to establish beam to T-944, but the MI-52 Septa would not turn on and would require a tunnel access to investigate further.
The Plans for Wednesday
The plans for today are to stack, stash, and store. MI is in access to repair the MI-52 Septa. The Run Coordinator will allow MI personnel to conduct studies.
Comments and Suggestions
|last modified 12/15/2010 email Fermilab|
|Security, Privacy, Legal|